OpenStudy (anonymous):

What is the solution set for |-4x - 8| = 4? Answer x = 3 and x = 1 x = 3 and x = -1 x = -3 and x = 1 x = -3 and x = -1 On the graph of the equation 3x - 2y = -18, what is the value of the y-intercept? Answer -9 -6 6 9

OpenStudy (anonymous):

Best to post one question at a time . . . For the first, consider that absolute value will make whatever is within positive, so you have to consider the positive case and the negative case.

OpenStudy (anonymous):

My apologies.....

OpenStudy (anonymous):

My answer to the first is D

OpenStudy (anonymous):

Easy way to do this is to split it into two separate equations: \[\large |-4x - 8| = 4 \rightarrow -4x-8 = 4; -4x-8 = -4\]

OpenStudy (anonymous):

{-3, -1} seem to work. It's a good habit to check the solutions with the original equation to make sure both are valid.

OpenStudy (anonymous):

For 3x - 2y = -18 This is a linear equation in standard form, Ax+By=C The x-intercept is at C/A, and the y-intercept is at C/B.

OpenStudy (anonymous):

I understand thus far